# Daedalus: Self-Adaptive Horizontal Autoscaling for Resource Efficiency of Distributed Stream Processing Systems

### 04 Mar 2024 v1

@@ 1. Introduction
Automated scaling of computational resources, commonly known as autoscaling, serves as the key method for aligning resources with dynamic workloads in DSP systems. The scaleout, defined by the number of worker nodes and processing slots per node, directly determines the system’s level of parallelism and influences the overall resource allocation. This, in turn, affects each worker node’s processing potential. Popular DSP frameworks like Apache Flink (Carbone et al., 2015), Kafka Streams (Sax et al., 2018), and Apache Spark (Zaharia et al., 2016), support dynamic adjustments in scaleout during runtime, which impacts the job’s parallelism and subsequently its resource utilization. Typically, these systems are deployed in cloud environments, where resources are provisioned elastically and can be scaled in or out as required, directly affecting the processing capabilities of the DSP job
@@ 4.1. DSP Jobs
Daedalus is evaluated with three representative DSP jobs: WordCount, Yahoo Streaming Benchmark, and Traffic Monitoring. All jobs, generators, and relevant code can be found in the Daedalus GitHub repository [^2] https://github.com/dos-group/daedalus.
@@ 4.3.3. Phoebe
Phoebe is an approach fairly similar to Daedalus, and unlike many other DSP autoscaling approaches, its source code is publicly available [^4] https://github.com/dos-group/phoebe, Accessed: March 2024. Phoebe uses initial profiling runs to build QoS models and TSF to better meet QoS requirements. Similarly to Daedalus, it is capable of choosing worker parallelism in accordance with a target recovery time. Unlike Daedalus though, Phoebe explicitly models and accounts for latency. It also injects failures into profiling runs, measures the resulting recovery times, and incorporates them into its QoS models. Phoebe was implemented using Flink and also evaluated using the Yahoo Streaming Benchmark. It will be compared to Daedalus using the target recovery time of 180 seconds, the same as in the original paper.
@@ 4.4. Experimental Setup
Approaches are evaluated by the quality of their scaling decisions. All jobs use exactly-once processing semantics and the DSP systems process all tuples. Therefore, end-to-end latency is a more suitable metric to indicate the processing performance. End-to-end latencies are reported by an empirical cumulative distribution function. This is a sorted latency measurement that describes the proportional percentage of observed latencies over the course of an experiment. Average latencies are also reported, though they can be misleading, since peaks of high latency during rescaling can skew an average upwards. While some papers exclude rescaling downtime in result calculations, this is misleading as recovery time is a crucial factor. Therefore, no metrics are excluded in the results. Lastly, the number of workers over the course of the experiment are shown and the normalized resource utilization is reported.
@@ 4.5.1. WordCount Results
The distribution of observed latencies is expressed as an empirical cumulative distribution function in 6(c). Generally, all approaches perform very similarly with most latency measurements falling between $10^{2}$ms and $10^{3}$ms. Average latencies over the span of the job are as follows: Daedalus with 1,171 ms, HPA 80 with 1,791 ms, HPA 85 with 961 ms, and Static 12 with 1,408 ms. Notably, the static scale-out has proportionally slightly higher latencies. As found in previous research (Zhang et al., 2021) and shown in 1(d), over-provisioning resources does not guarantee optimal latencies. Proportionally, the autoscaling approach latencies are quite similar. The larger increase in latencies above the 95% mark indicates when the systems are temporarily unavailable due to autoscaling.
@@ 4.6. Kafka Streams Experiment Results
Compared to the deployments that were able to process the workload, Daedalus used fewer resources. On average, Daedalus uses 5.2 workers, HPA 60 uses 5.8, HPA 80 4, and the static scale-out naturally used 12. Daedalus uses 57% less resources than the static scale-out and 11% less resources than the HPA 60. From these results, one can infer that Daedalus is a generally applicable solution. With Kafka Streams, it was able to provide a stable level of service using minimal resources.
@@ 4.7. Comparison with Phoebe
When comparing resources used during the autoscaling part of the experiment, Daedalus used 19% less resources, using an average of 10.1 workers while Phoebe used an average of 12.4 workers. However, Phoebe also requires initial profiling runs to build performance models. When incorporating profiling time, Daedalus used 53% less resources. Of course, the impact of this upfront cost decreases over the course of a long-running job. And so, while Phoebe is better able to target optimal latencies, it comes at the trade-off of increased resource usage.
@@ 4.8. Discussion
Since the quality of scaling decisions are also impacted by TSF, the accuracy of TSF predictions was evaluated. TSF predictions were generally accurate with errors typically falling below 5%. In fact, the threshold for poor predictions at 25% was never reached. The average WAPE was 2.2% for WordCount, 1.9% for the Yahoo Streaming Benchmark, and 2.0% for the Traffic Monitoring job.

### 05 Mar 2024 v2

@@ 1. Introduction
Automated scaling of computational resources, commonly known as autoscaling, serves as the key method for aligning resources with dynamic workloads in DSP systems. The scaleout, defined by the number of worker nodes and processing slots per node, directly determines the system’s level of parallelism and influences the overall resource allocation. This, in turn, affects each worker node’s processing potential. Popular DSP frameworks like Apache Flink (Carbone et al., 2015), Kafka Streams (Sax et al., 2018), and Apache Spark (Zaharia et al., 2016), support dynamic adjustments in scaleout during runtime, which impacts the job’s parallelism and subsequently its resource utilization. Typically, these systems are deployed in cloud environments, where resources are provisioned elastically and can be scaled in or out as required, directly affecting the processing capabilities of the DSP job.
@@ 3.4. Recovery Time
3.5. Monitoring with Anomaly Detection
To improve the accuracy of recovery time predictions, the actual recovery time after a scaling action is observed using statistical anomaly detection. When monitoring recovery time, the goal is to identify when processing returns back to normal after rescaling. Anomaly detection is therefore a fitting paradigm, as it provides a means for classifying normal and abnormal behavior. Because the objective is to find points when the system’s throughput deviates from the incoming workload, it is sufficient to use statistical anomaly detection on the difference between the workload and throughput. Statistical anomaly detection classifies observations as anomalous if the distance of an observation to the mean is above a certain threshold. Daedalus uses the threshold of one standard deviation. The anomaly detection model keeps track of the job’s running mean and variance of the difference between workload and throughput using Welford’s previously mentioned online algorithm. After a rescaling action, Daedalus checks for anomalies until the system has recovered. Because it takes time for a system to recover, the anomaly detection monitoring is run in a background thread to prevent interfering with the MAPE-K loop.
@@ 4.1. DSP Jobs
Daedalus is evaluated with three representative DSP jobs: WordCount, Yahoo Streaming Benchmark, and Traffic Monitoring. All relevant code can be found in the Daedalus GitHub repository [^2] https://github.com/dos-group/daedalus.
@@ 4.3.3. Phoebe
Phoebe is an approach fairly similar to Daedalus, and unlike many other DSP autoscaling approaches, its source code is publicly available [^4] https://github.com/dos-group/phoebe, Accessed: March 2024. Phoebe uses initial profiling runs to build QoS models and TSF to better meet QoS requirements. Similarly to Daedalus, it is capable of choosing worker parallelism in accordance with a target recovery time. Unlike Daedalus though, Phoebe explicitly models and accounts for latency. It also injects failures into profiling runs, measures the resulting recovery times, and incorporates them into its QoS models. Phoebe was implemented using Flink and also evaluated using the Yahoo Streaming Benchmark.
@@ 4.4. Experimental Setup
Approaches are evaluated by the quality of their scaling decisions. All jobs use exactly-once processing semantics and the DSP systems process all tuples. Therefore, end-to-end latency is a more suitable metric to indicate the processing performance. End-to-end latencies are reported by an empirical cumulative distribution function. Average latencies are also reported, though they can be misleading, since peaks of high latency during rescaling can skew an average upwards. While some papers exclude rescaling downtime in result calculations, this is misleading as recovery time is a crucial factor. Therefore, no metrics are excluded in the results.
@@ 4.5.1. WordCount Results
As can be seen in 6(c), generally, all approaches perform very similarly with most latency measurements falling between $10^{2}$ms and $10^{3}$ms. Average latencies over the span of the job are as follows: Daedalus with 1,171 ms, HPA 80 with 1,791 ms, HPA 85 with 961 ms, and Static 12 with 1,408 ms. Notably, the static scale-out has proportionally slightly higher latencies. As found in previous research (Zhang et al., 2021) and shown in 1(d), over-provisioning resources does not guarantee optimal latencies. Proportionally, the autoscaling approach latencies are quite similar. The larger increase in latencies above the 95% mark indicates when the systems are temporarily unavailable due to autoscaling.
@@ 4.6. Kafka Streams Experiment Results
Compared to the deployments that were able to process the workload, Daedalus used fewer resources. On average, Daedalus used 5.2 workers, HPA 60 used 5.8, HPA 80 used 4, and the static scale-out naturally used 12. Daedalus used 57% less resources than the static scale-out and 11% less resources than the HPA 60. From these results, one can infer that Daedalus is a generally applicable solution. With Kafka Streams, it was able to provide a stable level of service using minimal resources.
@@ 4.7. Comparison with Phoebe
When comparing resources used during the autoscaling part of the experiment, Daedalus used 19% less resources, using an average of 10.1 workers while Phoebe used an average of 12.4 workers. However, Phoebe also requires initial profiling runs to build performance models. When incorporating profiling time, Daedalus used 53% less resources.
@@ 4.8. Discussion
Since the quality of scaling decisions are also impacted by TSF, the accuracy of TSF predictions was evaluated. TSF predictions were generally accurate with errors typically falling below 5%. In fact, the threshold for poor predictions at 25% was never reached.